A firefighting crew uses a water cannon that shoots water at 25.0 m/s at a fixed angle of 47.0 ∘ above the horizontal. The firefighters want to direct the water at a blaze that is 15.0 m above ground level.
1
Expert's answer
2019-09-30T10:21:10-0400
"y=v_{0y}t-0.5gt^2=v_{0}\sin{47\degree}t-0.5gt^2"

"15=25\sin{47\degree}t-0.5(9.8)t^2"

"t_1=1.218\ s, t_2=2.513\ s"

"d_1=25\cos{47\degree}(1.218)=20.8\ m"

"d_2=25\cos{47\degree}(2.513)=42.9\ m"